A current, I=200A is passed through a copper conductor of cross section: S=1cm2. We suppose that all the atoms are transformed into Cu2+ ions. The density of copper is given by 8.9x103kg/m3 ; ( atomic mass is M=0.0635kg/mol and the Avogadro number is N=6.22x103 mol-1. Calculate the number of electrons per m3:



1
Expert's answer
2021-07-26T17:12:03-0400

Gives "I=200A\\S=1cm^2M=8.9\times10^3kg/m^3\\M=0.0635kg/mol\\N=6.23\times10^{23}mol^{-1}\\cu^{2+},n=2"

"\frac{N}{N_A}=\frac{m}{M}"

"n=\frac{N}{V}"

"n=\frac{m}{M}\times\frac{N_A}{V}"

"n=\frac{N_A}{V}\times d"

Electron density


"n=\frac{6.22\times10^{23}\times8.9\times10^3}{0.0635}=8.72\times10^{28}m^{-3}"

Now as such atom contribution two electron the number of electron per unit volume

"n_e=2\times n"


"n_e=2\times8.72\times10^{28}=17.4\times10^{28}m^{-3}"